OpenStudy (aravindg):

\lim_{x \rightarrow 0} e ^{(1/x*\log x)}

OpenStudy (jamesj):

\[ \lim_{x \rightarrow 0+} 1/x = +\infty \ \ and \ \ \lim_{x \rightarrow 0+} ln x = -\infty\] Hence \[ \lim_{x \rightarrow 0+} (ln x)/x = -\infty \] Thus ... what does this tell you about your limit?

OpenStudy (aravindg):

not defined??

OpenStudy (jamesj):

Oh, come on. What is \[\lim_{x \rightarrow -\infty} \ e^x\] ?

OpenStudy (aravindg):

my question is \[\lim_{x \rightarrow 0}e ^{(1/x*\log x)}\]

OpenStudy (aravindg):

i meant e raised to

OpenStudy (jamesj):

Yes, I just showed you how the index goes to -infinity. Now what happens to the function e^x as x --> -infinity?

OpenStudy (aravindg):

not defined is the answer?

OpenStudy (jamesj):

No. Let me brak this down ever further for you. Let y(x) = (ln x)/x Then limit y(x) as x --> 0+ is -infinity, agreed?

OpenStudy (jamesj):

Now what is \[\lim_{y \rightarrow -\infty} e^y\]

OpenStudy (aravindg):

yes

OpenStudy (aravindg):

-infinity

OpenStudy (jamesj):

No ... that's a bad error. What does the graph of the exponential function look like?

OpenStudy (aravindg):

pls tel final answer i hav many othr questions to do :(

OpenStudy (jamesj):

The limit of e^y as y --> -infinity is zero.

OpenStudy (jamesj):

hence your limit is what ... ?

OpenStudy (aravindg):

+infinity

OpenStudy (jamesj):

What is \[\lim_{y \rightarrow -\infty} \ e^y\] ?

OpenStudy (turingtest):

sorry to butt in, but please James, just have a look at my question when you get the chance, it should be easy for you.

OpenStudy (aravindg):

tl the FINAL ANSWER PLZZZZZZZZZZZZZZZZZZZZZZZ

OpenStudy (jamesj):

@TT: ok

OpenStudy (jamesj):

If you don't understand this basic property of the exponential that the limit as y-->-infinity is zero, you'll be in deep trouble later. The answer is this: \[y(x) = (\ln x) / x\] \[\lim_{x \rightarrow 0+} y(x) = -\infty \] \[ \lim_{y \rightarrow -\infty} e^y = 0\] therefore \[\lim_{x \rightarrow 0+} e^{(ln x)/x} = 0\]